the space shuttle is on its way to Mars. The inside of the shuttle is pressurized to 1 atm. The windshield is a rectangle 2.5m by 1m. How much force is applied to the windshield?
Answer #1

Force =Pressure * area =101325 *2.5 *1=253312.5 N
Answer #2

Force(F) = Pressure(P) x Area(A)

1 atm = 1.01325*10^5 Pa

Area = 2.5*1 = 2.5 m^2

F =  1.01325*10^5*2.5 = 2.533125*10^5 N

F = 2.533125*10^5 N
Answer #4

FORCE = PRESSURE * AREA

=101325 * 2.5

=253312.5
Answer #5

In the space, pressure outside of shuttle = 0 N/m2

Inside Pressure = 1 atm = 1.01325